OpenStudy (anonymous):

If you know that is a root of a polynomial function, then the name given to 3 minus the square root of 11, another root of that same function, is a _____ conjugate.

OpenStudy (anonymous):

\(3+\sqrt{11}\) is the conjugate

OpenStudy (anonymous):

conjugate of \(a+\sqrt{b}\) is \(a-\sqrt{b}\)